# A mesh-constrained discrete point method for incompressible flows with moving boundaries

### v1

@@ Abstract
Particle-based methods are a practical tool in computational fluid dynamics, and novel types of methods have been proposed. However, widely developed Lagrangian-type formulations suffer from the nonuniform distribution of particles, which is enhanced over time and result in problems in computational efficiency and parallel computations. To mitigate these problems, a mesh-constrained discrete point (MCD) method was developed for stationary boundary problems (Matsuda et al., 2022). Although the MCD method is a meshless method that uses moving least-squares approximation, the arrangement of particles (or discrete points (DPs)) is specialized so that their positions are constrained in background meshes to obtain a closely uniform distribution. This achieves a reasonable approximation for spatial derivatives with compact stencils without encountering any ill-posed condition and leads to good performance in terms of computational efficiency. In this study, a novel meshless method based on the MCD method for incompressible flows with moving boundaries is proposed. To ensure the mesh constraint of each DP in moving boundary problems, a novel updating algorithm for the DP arrangement is developed so that the position of DPs is not only rearranged but the DPs are also reassigned the role of being on the boundary or not. The proposed method achieved reasonable results in numerical experiments for well-known moving boundary problems.
@@ 1 Introduction
The uniformity of the particle distribution is an important feature for achieving stable and efficient computations in particle methods because the local arrangement of particles influences conditions in the evaluation of spatial gradients. Particularly, in pure Lagrangian particle methods, the heterogeneous distribution of particles often occurs according to their motion [20], and exhibits sparsity and clustering because of tensile instability [21, 22, 23]. This reduces numerical stability and accuracy. In this regard, a large number of stencils is required for the stable calculation of derivative values; however, this results in a large computational cost compared with mesh-based methods.
@@
The ALE formulation also has the potential to reduce the bottleneck for load-in balance in parallel computation by controlling particle positions adequately. Despite this, even when efficient parallelization techniques [36, 37, 38] are used, it is still challenging to establish a universal (i.e., problem-independent) method that achieves a good balance for calculation and node-to-node communication because neighboring information about the particle arrangement should be reassigned according to particle motions.
The mesh-constrained discrete point (MCD) method was designed to mitigate aforementioned problems [39]. Although the MCD method is a meshless method that uses MLS approximation, the arrangement of particles (or DPs) is specialized so that their positions are constrained in background meshes to obtain a closely uniform distribution. This achieves a reasonable approximation for spatial derivatives with compact stencils without encountering any ill-posed condition and leads to good performance in terms of computational efficiency. Unlike existing hybrid methods that use particle and mesh systems, such as the material point method ([40, 41, 42, 43, 44]) and particle-grid hybrid method ([45, 46, 47, 48, 49]), the MCD method only uses background meshes to determine the DP positions, and does not need any interpolation between particles and mesh systems, which appears in other hybrid methods. At present, the MCD method is limited to unsteady Stokes flows with stationary boundaries, and a novel formulation needs to be applied for moving boundary flow problems.
In this study, a novel meshless method is proposed based on the MCD method for incompressible flows with moving boundaries. To ensure the mesh constraint of each DP in moving boundary problems, a novel updating algorithm for the DP arrangement is developed so that the position of DPs is not only rearranged but the DPs are also reassigned the role of being on the boundary or not. Because each DP is prohibited from moving outside the background mesh, capturing a moving interface using the same DPs is not possible. This problem is solved by changing the DP roles (domain/boundary/void) according to information about the relative positions between background meshes and moving boundaries, and controlling the DP arrangement along interface motions. The proposed method achieved reasonable results in numerical experiments for well-known moving boundary problems.
@@ 2 MCD method
2.1 Representation of arbitrary boundary shapes with a signed distance function (SDF)
@@ 2.2 DP masking algorithm
In the MCD method, to identify meshes and DPs for computation with moving boundaries, “mask” is introduced. In [39], temporal points (TPs) were introduced on the centroids of background meshes to determine the mask according to the signed distance between the TP and boundary. In the present study, an idea is developed to assign the mask based on the positional relationship between mesh vertices and moving boundaries, rather than using TPs. In this study, a single DP is set in each mesh, that is, the common mask is assigned to the DP and corresponding mesh. Let $n_{i,\,j}^{\rm mesh}$ denote the number of vertices, at where the SDF is greater than or equal to zero, for a mesh ($i,~{}j$). In a 2D uniform Cartesian mesh system with mesh length $l_{0}$, $n_{i,\,j}^{\rm mesh}$ can take values from 0 to 4 because the mesh ($i,~{}j$) has four vertices: ($i,~{}j$), ($i+1,~{}j$), ($i,~{}j+1$), and ($i+1,~{}j+1$). The mask at ($i,j$) is defined as
${\rm mask}_{i,\,j}=\left\{\begin{array}[]{ll}0,&\text{for}~{}n_{i,\,j}^{\rm mesh}=0,~{}1,\\ 1,~{}2,~{}...,&\text{for}~{}n_{i,\,j}^{\rm mesh}=2,~{}3,\\ -1,~{}-2,~{}...,&\text{for}~{}n_{i,\,j}^{\rm mesh}=4,\end{array}\right.$ (1)
where the value indicates that ${\rm mask}_{i,\,j}\geq 1$ for (moving) boundaries; ${\rm mask}_{i,\,j}\leq-1$ for the inner domain; and ${\rm mask}_{i,\,j}=0$ for the void. Note that different values for ${\rm mask}_{i,\,j}\geq 1$ can be used to identify multiple boundaries.
In this study, once the aforementioned masking is performed, one more operation is performed as follows: If mesh $(i,\,j)$ assigning the inner domain has left, right, upper, and lower neighboring meshes assigning the void, the mask of mesh $(i,\,j)$ is reassigned as the boundary. This operation is necessary to avoid a lack of DPs representing boundaries because of a singular arrangement between the background meshes and boundaries, for example, a linear boundary overlaps the mesh edge.
@@ 2.3 DP arrangement algorithm
    Set an initial DP arrangement at the centroids of background meshes $\mathbf{x}_{\rm CM}$ ($\mathbf{x}^{0}=\mathbf{x}_{\rm CM}$).
@@
    Move the DPs homogeneously by solving a mechanical equilibrium problem ($\mathbf{x}^{k}\rightarrow\mathbf{x}^{*}$).
@@
where $\kappa=\Delta\tau F_{0}/\gamma l_{0}$ denotes the dimensionless coefficient. When $\kappa$ is set to a certain value, the proposed algorithm exhibits similar convergence behavior [39]. In this study, $\kappa$ is set to $\kappa=1/5$.
@@
In procedure (iv), the mesh constraint is applied for DPs that exceed each background mesh, which are relocated to the nearest edge or vertex of the background mesh (Figure 3).
Figure 3: Examples of the mesh constraint procedure for a DP that exceeds the background mesh. (a) When the DP (dashed line) is outside of the mesh edge for the $x$-axis, its $x$-coordinate is replaced by that of the mesh edge (black dot). (b) Same as (a), but for the $y$-axis. (c) When the DP (dashed circle) is outside of both mesh edges for both $x$ and $y$-axes, its $x$ and $y$-coordinates are replaced by those of the mesh edges, which relocates the DP to the nearby mesh vertex (black dot).
@@ 2.4 Treatment of moving boundaries
Compared with fully Lagrangian particle methods, the main issue in applying the proposed MCD method to moving boundary problems is how to manage the moving boundary using the mesh-constrained approach. Figure 4 shows a conceptual illustration of the proposed approach to update the DP arrangement for the moving boundary. The DPs are masked according to the relationship between the arrangements of the background mesh and the moving boundary. A moving boundary that moves toward the left (Figure 4-a) is considered. If the relationships between the background mesh and moving boundary are not changed, the mask of each DP is maintained (Figure 4-b). When the boundary moves beyond the mesh vertices, the masks of DPs are updated according to a new arrangement (Figure 4-c). As in the enlarged part of the figure, the mask of the DP that was on $\Gamma$ at the previous time changed from “boundary” to “void” and the DP was excluded from the calculation, whereas the mask of the DP that was in $\Omega_{\rm I}$ at the previous time changed from “void” to “boundary” and the position was relocated onto $\Gamma$. A similar procedure was adopted for the DPs located on the opposite side of the direction of travel, that is, the mask of the DPs changed from “boundary” to “void” or “inner” to “boundary.”
Figure 4: Example of the mesh-constrained approach for a moving boundary. (a) DP arrangement around a moving boundary $\Gamma$ that moves toward the left. (b) Unless the relationship between the background mesh and moving boundary changes, the mask of each DP is maintained. (c) If a boundary moves beyond a mesh vertex, the mask and position of the DP is updated according to the new arrangement.
@@ 3.1 Governing equations
$\frac{\partial\mathbf{v}}{\partial t}+\left(\mathbf{v}-\mathbf{w}\right)\cdot\nabla\mathbf{v}=-\nabla P+\nu\nabla^{2}\mathbf{v},$ (11)
where $\mathbf{v}$, $P$, and $\nu$ denote the velocity vector, pressure normalized by the fluid density, and kinematic viscosity, respectively. Additionally, $t$ denotes time, $\nabla=\partial/\partial\mathbf{x}$ denotes the spatial differential operator, and $\mathbf{w}$ denotes the velocity of DPs (ALE velocity).
@@ 3.2 MLS approximation for spatial derivatives
Now the evaluation of the discrete quantity $\phi_{c}$ at an arbitrary position $\mathbf{x}_{c}$ is considered. In the MCD method, DPs can be distributed in arbitrary positions around boundary shapes under the mesh constraint condition. Hence, in this study, the generalized quadratic MLS formulation described in [39] is used for evaluating a discretized quantity and its spatial differentiation. The MCD method can use the structure information for adjacent DPs using background mesh systems and execute stencil calculations efficiently, which is a different feature from other particle-based methods that have no DP adjacent information. In this study, 2D systems are calculated using a compact support domain $D_{c}$ of $3\times 3$ background meshes for arbitrary point $\mathbf{x}_{c}$.
@@
3.3 Velocity-pressure coupling
@@
In this study, the series is expanded to the second-order ($p=2$), which is equivalent to interpolation using a quadratic MLS reconstruction for $\chi_{c}=1$ at ${\bf x}^{k}$. Note that, in the framework of the proposed method, the DP not only moves along the moving boundary but also changes its mask. The case in which the DP on the boundary in the $k$-th step is updated is considered, with its mask as the domain, and relocated in the next step. In this case, the physical information (e.g., pressure or velocity) of the DP after its mask is changed is missing. To uniformly obtain $\tilde{\mathbf{u}}^{k}$ of the DP cloud in $\Omega_{\rm I}$, including the DP that is missing information, the previous (or $k$-th) mask, position, and velocity are used as the stencil calculation for the MLS interpolation (Appendix A.).
@@
$\frac{\partial P}{\partial n}=-\frac{{\rm d}\mathbf{V}}{{\rm d}t}^{k+1}\cdot\mathbf{n}^{k+1},$ (23)
@@
where $\bm{\omega}$ denotes the angular velocity and $\mathbf{r}$ denotes the direction vector from the centroid. The temporal derivative of $\mathbf{V}$ at $k+1$ is approximated by
$\frac{{\rm d}\mathbf{V}}{{\rm d}t}^{k+1}$ $=\frac{{\rm d}\mathbf{V}_{\rm trans}}{{\rm d}t}^{k+1}+\frac{{\rm d}\bm{\omega}}{{\rm d}t}^{k+1}\times\mathbf{r}^{k+1}+\bm{\omega}^{k+1}\times\mathbf{V}^{k+1}$ (26)
$\approx\frac{\mathbf{V}_{\rm trans}^{k+1}-\mathbf{V}_{\rm trans}^{k}}{\Delta t}+\frac{\bm{\omega}^{k+1}-\bm{\omega}^{k}}{\Delta t}\times\mathbf{r}^{k+1}+\bm{\omega}^{k+1}\times\mathbf{V}^{k+1}.$ (27)

### v3

@@ Abstract
Particle-based methods are a practical tool in computational fluid dynamics, and novel types of methods have been proposed. However, widely developed Lagrangian-type formulations suffer from the nonuniform distribution of particles, which is enhanced over time and result in problems in computational efficiency and parallel computations. To mitigate these problems, a mesh-constrained discrete point (MCD) method was developed for stationary boundary problems (Matsuda et al., 2022). Although the MCD method is a meshless method that uses moving least-squares approximation, the arrangement of particles (or discrete points (DPs)) is specialized so that their positions are constrained in a background mesh to obtain a closely uniform distribution. This achieves a reasonable approximation for spatial derivatives with compact stencils without encountering any ill-posed condition and leads to good performance in terms of computational efficiency. In this study, a novel meshless method based on the MCD method for incompressible flows with moving boundaries is proposed. To ensure the mesh constraint of each DP in moving boundary problems, a novel updating algorithm for the DP arrangement is developed so that the position of DPs is not only rearranged but the DPs are also reassigned the role of being on the boundary or not. The proposed method achieved reasonable results in numerical experiments for well-known moving boundary problems.
@@ 1 Introduction
The uniformity of the particle distribution is an important feature for achieving stable and efficient computations in particle methods because the local arrangement of particles influences conditions in the evaluation of spatial gradients. Particularly, in pure Lagrangian particle methods, the nonuniform distribution of particles often occurs according to their motion [20], and exhibits sparsity and clustering because of tensile instability [21, 22, 23]. This reduces numerical stability and accuracy. In this regard, a wide stencil is required for the stable calculation of derivative values; however, this results in a large computational cost compared with mesh-based methods.
@@
The ALE formulation also has the potential to reduce the bottleneck for load imbalance in parallel computation by controlling particle positions adequately. Despite this, even when efficient parallelization techniques [36, 37, 38] are used, it is still challenging to establish a universal (i.e., problem-independent) method that achieves a good balance for calculation and node-to-node communication because neighboring information about the particle arrangement should be reassigned according to particle motions.
The mesh-constrained discrete point (MCD) method was designed to mitigate aforementioned problems [39]. Although the MCD method is a meshless method that uses MLS approximation, the arrangement of particles (or DPs) is specialized so that their positions are constrained in a background mesh to obtain a closely uniform distribution. This achieves a reasonable approximation for spatial derivatives with compact stencils without encountering any ill-posed condition and leads to good performance in terms of computational efficiency. Unlike existing hybrid methods that use particle and mesh systems, such as the material point method ([40, 41, 42, 43, 44]) and particle-grid hybrid method ([45, 46, 47, 48, 49]), the MCD method only uses the background mesh to determine the DP positions, and does not need any interpolation between particles and mesh systems, which appears in other hybrid methods. At present, the MCD method is limited to unsteady Stokes flows with stationary boundaries, and a novel formulation needs to be applied for moving boundary flow problems.
In this study, a novel meshless method is proposed based on the MCD method for incompressible flows with moving boundaries. To ensure the mesh constraint of each DP in moving boundary problems, a novel updating algorithm for the DP arrangement is developed so that the position of DPs is not only rearranged but the DPs are also reassigned the role of being on the boundary or not. Because each DP is prohibited from moving outside the background mesh element, capturing a moving interface using the same DPs is not possible. This problem is solved by changing the DP roles (domain/boundary/void) according to information about the relative positions between background mesh elements and moving boundaries, and controlling the DP arrangement along interface motions. The proposed method achieved reasonable results in numerical experiments for well-known moving boundary problems.
@@ 2 MCD method
2.1 Overview
In the MCD method, the DPs for solution unknowns are dispersedly arranged in the fluid domain and (moving) boundary that allows to represent arbitrary boundary shapes accurately. A unique feature of the MCD method is to introduce a background mesh system to constrain the DP arrangement, where each DP is assigned to a mesh element and prohibited from moving over corresponding mesh edges. A mechanical-based relaxation algorithm is applied to maintain smooth distribution of DPs and strict restriction algorithm is applied for satisfying the mesh constraint. Although a main idea has been established in [39], in this study, little modification for the DP masking and arrangement is introduced to simplify the algorithm. To manage DPs with moving boundaries, a DP relocation algorithm is applied in combination with the mesh-constrained approach. In this study, an arbitrary shape is given by a discrete signed distance function (SDF), that simplifies to treat a non-deformable moving body and to combine with the DP masking and arrangement algorithms.
@@ 2.2 DP masking algorithm
In the MCD method, to identify mesh elements and DPs for computation with moving boundaries, “mask” is introduced. In [39], temporal points (TPs) were introduced on the centroids of each background mesh element to determine the mask according to the SDF between the TP and boundary. In the present study, an idea is developed to assign the mask based on the positional relationship between mesh vertices and moving boundaries, rather than using TPs. A single DP is set in each mesh element, that is, the same mask is assigned to both the DP and mesh. Let $n_{i,\,j}^{\rm mesh}$ denote the number of vertices, at where the SDF is greater than or equal to zero, for a mesh ($i,~{}j$). In a 2D uniform Cartesian mesh system with mesh length $l_{0}$, $n_{i,\,j}^{\rm mesh}$ can take values from 0 to 4 because the mesh ($i,~{}j$) has four vertices: ($i,~{}j$), ($i+1,~{}j$), ($i,~{}j+1$), and ($i+1,~{}j+1$). The mask at ($i,j$) is defined as
${\rm mask}_{i,\,j}=\left\{\begin{array}[]{ll}0,&\text{for}~{}n_{i,\,j}^{\rm mesh}=0,~{}1,\\ 1,~{}2,~{}...,&\text{for}~{}n_{i,\,j}^{\rm mesh}=2,~{}3,\\ -1,&\text{for}~{}n_{i,\,j}^{\rm mesh}=4,\end{array}\right.$ (1)
where the value indicates that ${\rm mask}_{i,\,j}\geq 1$ for (moving) boundaries; ${\rm mask}_{i,\,j}=-1$ for the inner domain; and ${\rm mask}_{i,\,j}=0$ for the void. Note that different values for ${\rm mask}_{i,\,j}\geq 1$ can be used to identify multiple boundaries.
@@ 2.3 DP arrangement algorithm
    Set an initial DP arrangement at the centroid $\mathbf{x}_{\rm CM}$ of each background mesh element ($\mathbf{x}^{0}=\mathbf{x}_{\rm CM}$).
@@
    Move the DPs uniformly in space by solving a mechanical equilibrium problem ($\mathbf{x}^{k}\rightarrow\mathbf{x}^{*}$).
@@
where $\kappa=\Delta\tau F_{0}/\gamma l_{0}$ denotes the dimensionless coefficient. When $\kappa$ is set to a certain value, the proposed algorithm exhibits similar convergence behavior regardless of the mesh length [39]. In this study, $\kappa$ is set to $\kappa=1/5$.
@@
In procedure (iv), the mesh constraint is applied for DPs that exceed each background mesh element, which are relocated to the nearest edge or vertex of the background mesh element (Figure 2).
Figure 2: Examples of the mesh constraint procedure for a DP that exceeds the element of background mesh corresponding to the DP. (a) When the DP (dashed line) is outside of the edges of the mesh element for the $x$-axis, its $x$-coordinate is replaced by that of the edges of the mesh element (black dot). (b) Same as (a), but for the $y$-axis. (c) When the DP (dashed circle) is outside of both edges of the mesh element for both $x$ and $y$-axes, its $x$ and $y$-coordinates are replaced by those of the edges of the mesh element, which relocates the DP to the nearby vertex of the mesh element (black dot).
@@ 2.4 Treatment of moving boundaries
Compared with fully Lagrangian particle methods, the main issue in applying the proposed MCD method to moving boundary problems is how to manage the moving boundary using the mesh-constrained approach. Figure 3 shows a conceptual illustration of the proposed approach to update the DP arrangement for the moving boundary. The DPs are masked according to the positional relationship between each element of the background mesh and the moving boundary. A moving boundary that moves toward the left (Figure 3-a) is considered. If the relationships between each background mesh element and moving boundary are not changed, the mask of each DP is maintained (Figure 3-b). When the boundary moves across the mesh vertices, the masks of DPs are updated according to a new arrangement (Figure 3-c). As in the enlarged part of the figure, the mask of the DP that was on $\Gamma$ at the previous time changed from “boundary” to “void” and the DP was excluded from the calculation, whereas the mask of the DP that was in $\Omega_{\rm I}$ at the previous time changed from “void” to “boundary” and the position was relocated onto $\Gamma$. A similar procedure was adopted for the DPs located on the opposite side of the direction of travel, that is, the mask of the DPs changed from “boundary” to “void” or “inner” to “boundary.”
Figure 3: Example of the mesh-constrained approach for a moving boundary. (a) DP arrangement around a moving boundary $\Gamma$ that moves toward the left. (b) Unless the relationship between the background mesh elements and moving boundary changes, the mask of each DP is maintained. (c) If a boundary moves beyond a mesh vertex, the mask and position of the DP is updated according to the new arrangement.
@@ 3.1 Governing equations
$\frac{\partial\mathbf{v}}{\partial t}+\left(\mathbf{v}-\mathbf{w}\right)\cdot\nabla\mathbf{v}=-\nabla P+\nu\nabla^{2}\mathbf{v}.$ (11)
Here, $\mathbf{v}$ is the velocity vector, $P$ is the pressure normalized by the fluid density $\rho$, and $\nu=\mu/\rho$ is the kinematic viscosity, where $\mu$ is the dynamics viscosity. Additionally, $t$ denotes time, $\nabla=\partial/\partial\mathbf{x}$ denotes the spatial differential operator, and $\mathbf{w}$ denotes the velocity of DPs (ALE velocity).
@@ 3.2 MLS approximation for spatial derivatives
Now the evaluation of the discrete quantity $\phi_{c}$ at an arbitrary position $\mathbf{x}_{c}$ is considered. In the MCD method, DPs can be distributed in arbitrary positions around boundary shapes under the mesh constraint condition. Hence, in this study, the generalized quadratic MLS formulation described in [39] is used for evaluating a discretized quantity and its spatial differentiation. The MCD method can use the structure information for adjacent DPs using the background mesh system and execute stencil calculations efficiently, which is a different feature from other particle-based methods that have no DP adjacent information. In this study, 2D systems are calculated using a compact support domain $D_{c}$ of $3\times 3$ background mesh elements for arbitrary point $\mathbf{x}_{c}$.
@@
3.3 Velocity–pressure coupling
@@
In this study, the series is expanded to the second-order ($p=2$), which is equivalent to interpolation using a quadratic MLS reconstruction for $\chi_{c}=1$ at ${\bf x}^{k}$. Note that, in the framework of the proposed method, the DP not only moves along the moving boundary but also changes its mask. The case is considered that the DP on the boundary at $k$-th step becomes the one on the fluid domain at the $k+1$-th step. In this case, the physical information (e.g., pressure or velocity) of the DP after its mask is changed is missing. To uniformly obtain $\tilde{\mathbf{u}}^{k}$ of the DP cloud in $\Omega_{\rm I}$, including the DP that is missing information, the previous (or $k$-th) mask, position, and velocity are used as the stencil calculation for the MLS interpolation (Appendix A.).
@@
$\frac{\partial P}{\partial n}=-\frac{{\rm d}\mathbf{V}^{k+1}}{{\rm d}t}\cdot\mathbf{n}^{k+1},$ (23)
@@
where $\bm{\omega}$ denotes the angular velocity and $\mathbf{r}~{}(=\mathbf{x}_{\Gamma}-\mathbf{x}_{\rm c})$ denotes the direction vector from the centroid. The temporal derivative of $\mathbf{V}$ at $k+1$ is approximated by
$\frac{{\rm d}\mathbf{V}^{k+1}}{{\rm d}t}$ $=\frac{{\rm d}\mathbf{V}_{\rm trans}^{k+1}}{{\rm d}t}+\frac{{\rm d}\bm{\omega}^{k+1}}{{\rm d}t}\times\mathbf{r}^{k+1}+\bm{\omega}^{k+1}\times\frac{{\rm d}\mathbf{r}^{k+1}}{{\rm d}t}$ (26)
$=\frac{{\rm d}\mathbf{V}_{\rm trans}^{k+1}}{{\rm d}t}+\frac{{\rm d}\bm{\omega}^{k+1}}{{\rm d}t}\times\mathbf{r}^{k+1}+\bm{\omega}^{k+1}\times\left(\frac{{\rm d}\mathbf{x}_{\Gamma}^{k+1}}{{\rm d}t}-\frac{{\rm d}\mathbf{x}_{\rm c}^{k+1}}{{\rm d}t}\right)$ (27)
$=\frac{{\rm d}\mathbf{V}_{\rm trans}^{k+1}}{{\rm d}t}+\frac{{\rm d}\bm{\omega}^{k+1}}{{\rm d}t}\times\mathbf{r}^{k+1}+\bm{\omega}^{k+1}\times\mathbf{V}_{\rm rot}^{k+1}$ (28)
$\approx\frac{\mathbf{V}_{\rm trans}^{k+1}-\mathbf{V}_{\rm trans}^{k}}{\Delta t}+\frac{\bm{\omega}^{k+1}-\bm{\omega}^{k}}{\Delta t}\times\mathbf{r}^{k+1}+\bm{\omega}^{k+1}\times\mathbf{V}_{\rm rot}^{k+1},$ (29)
where, ${\rm d}\mathbf{x}_{\Gamma}/{\rm d}t=\mathbf{V}_{\rm trans}+\mathbf{V}_{\rm rot}$ and ${\rm d}\mathbf{x}_{c}/{\rm d}t=\mathbf{V}_{\rm trans}$.